Sala Co. is contemplating the replacement of an old machine with a new one. The following information has been gathered: Old Machine New Machine Price $250,000 $500,000 Accumulated Depreciation 75,000 -0- Remaining useful life 10 years -0- Useful life -0- 10 years Annual operating costs $200,000 $150,500 If the old machine is replaced, it can be sold for $20,000. The net advantage (disadvantage) of replacing the old machine is:

a. $15000

b.$20000

c.(5,000)

d.(50,000)
Answer #1
Old New
Operating expenses 200,000 150,000
Depreciation 17500 50000
Total expenses 217,500 200,000
Decrease in expenses due to new machine
[217500- 200000]
17500
Decrease in the expenses for 10 years 175000
Less: Loss of disposal of machine
[250000-75000-20000]
155000
Net advantage 20000

The correct option is (b).

Depreciation for old machine = (250,000-75,000)/10 = 17,500

Depreciation for new machine = 500,000/10 = 50,000